Please show me the steps for finding an inverse for
y= 3x-6
I have to solve for y
also, Y= x^2 + 5

Thank you so much.

Answer by lagbens(14)

to solve this you need to replace x by y and y by x and then solve for y.
so y=3x-6 would turn into x=3y-6.
okay now let's solve for y; x=3y-6 (add 6 to both side)
x+6=3y(divide both side by 3)
y=x+6/3 (final answer)
follow the same steps for the next one. hope this helps.